OpenStudy (anonymous):

Using radicals as needed, multiply (sqrt7 +1)(8sqrt7 +1)

OpenStudy (anonymous):

\[(\sqrt7 +1) (8\sqrt7 +1)\]

jimthompson5910 (jim_thompson5910):

\[(\sqrt{7}+1)(8\sqrt{7} +1) \] \[(\sqrt{7})(8\sqrt{7})+(\sqrt{7})(1)+(1)(8\sqrt{7})+1*1 \] \[8\sqrt{7*7}+\sqrt{7}+8\sqrt{7}+1\] \[8\sqrt{49}+\sqrt{7}+8\sqrt{7}+1\] \[8(7)+\sqrt{7}+8\sqrt{7}+1\] \[56+\sqrt{7}+8\sqrt{7}+1\] \[57+9\sqrt{7}\] So \[(\sqrt{7}+1)(8\sqrt{7} +1) = 57+9\sqrt{7} \]